the sun is 1.50 x 108 km from earth, and the speed of light is 3.00 x 108 m/s. how many minutes elapse as light travels from the sun toearth?

please show work
Answer #1

1.5 x 10^8 x 1000=1.5 x 10^11meter     s=d/t     3.00 x 10^8=1.5 x 10^11/t  

t=1.5 x 10^11/3.00 x 10^8=1.5 x 10^3 / 3.00=1 x 10^3/2=1000/2=500 seconds

t=500/60=8.33 minutes

we should convert km to meter and seconds to minutes